OpenStudy (anonymous):

The equation of a circle is (x + 4)2 + (y + 2)2 = 16. Determine the coordinates of the center of the circle and the length of the radius.

OpenStudy (anonymous):

Answers Center (-2, -4), radius is 16 Center (2, 4), radius is 16 Center (4, 2), radius is 4 Center (-4, -2), radius is 4

OpenStudy (anonymous):

@Loser66 @AlexandervonHumboldt2 @bibby @CausticSyndicalist @Directrix @dtan5457 @Mehek14 @mathmath333

OpenStudy (loser66):

What is the equation of a circle?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

ok so the equation of a circle is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

and so in your equation you have a +4 and +2 and your radius squared is 16. what is the square root of 16? also in the equation of a circle the h and k are negative but in your problem they are positive why do you think that is?

OpenStudy (anonymous):

@highjumpcollect

OpenStudy (anonymous):

sorry i had to use the bathroom

OpenStudy (anonymous):

it's ok :)

OpenStudy (anonymous):

SR of 16=4 would it be third choice?

OpenStudy (anonymous):

wait what is the standard equation of a circle?

OpenStudy (anonymous):

so youo eliminated the first two choices but look at the circle equation and your equation in the circle equation the h and k are negative but in your equation they are positive how do you turn a negative into a positive?

OpenStudy (anonymous):

the standard equation is (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

i have no idea.. this question isn't my favorite at all! lol

OpenStudy (anonymous):

lol trusst me i had trouble learning the circle stuff too wait till you get to the unit circle..... but a negative times a negative equals a positive

OpenStudy (anonymous):

duhh ok right i remember that.. its so simple its hard

OpenStudy (anonymous):

so it would e the 3rd choice? or 4th?

OpenStudy (anonymous):

4th

OpenStudy (anonymous):

Thank u SO much

OpenStudy (anonymous):

lol you're welcome